# revtex4-2Repair the float

### v1

@@ III.2 Spin-triplet superconductor junctions with the time-reversal symmetry
Next, we demonstrate the tunneling conductance of the spin-triplet SC junctions with the time-reversal symmetry. We plot $\sigma(eV,k_{x})$, $\sigma_{\uparrow}(eV,k_{x})$, and $\sigma_{\downarrow}(eV,k_{x})$ for helical $p$-wave, $p_{x}$-wave, and $p_{y}$-wave SC junctions in Figs. 5 (a)(d)(g), Figs. 5 (b)(e)(h), and Figs. 5 (c)(f)(i), respectively. In the case of the helical $p$-wave junction, we obtain the missing of the helical edge states from $-k_{x2}\leq k_{x}\leq-k_{x1}$ and $k_{x1}\leq k_{x}\leq k_{x2}$ as seen from $\sigma(eV,k_{x})$ plotted in Fig. 5 (a). For the $p_{x}$-wave SC junctions, because we consider the junction along the $y$-direction, we obtain the nodal structure in the conductance [Fig. 5 (b)] and ZBCP is absent. On the other hand, in the $p_{y}$-wave SC junctions, because we consider the junction along the $y$-direction, we obtain ZBCP caused by the zero-energy ABSs of the $p_{y}$-wave pairing [Fig. 5 (c)]. When we see the tunneling conductance with up-spin $\sigma_{\uparrow}(eV,k_{x})/\sigma_{\mathrm{N}}$, the range of $k_{x}$ contributing to the conductance is restricted for $-k_{x2}\leq k_{x}\leq k_{x1}$ [Figs. 5 (d)(e)(f)]. On the other hand, for tunneling conductance with down spin, $\sigma_{\downarrow}(eV,k_{x})/\sigma_{\mathrm{N}}$ becomes nonzero for $-k_{x1}\leq k_{x}\leq k_{x2}$ [Figs. 5 (g)(h)(i)]. We also find the relation: $\sigma_{\uparrow}(eV,k_{x})=\sigma_{\downarrow}(eV,-k_{x})$.
The resulting tunneling conductance $\bar{\sigma}(eV)$ is plotted in Fig. 6 with $t_{J}=0$ and $t_{J}=0.25t$ shown in Figs. 6 (a)(b)(c) and Figs. 6 (d)(e)(f). For the helical $p$-wave SC junctions, in the absence of the PUM order, the tunneling conductance is symmetric for $eV$ [Fig. 6 (a)]. However, because the momentum resolved tunneling conductance $\sigma(eV,k_{x})$ is asymmetric for $eV$ [Fig. 5 (a)], $\bar{\sigma}(eV)$ shown in Fig. 6 (f) is not symmetric for $eV$. This asymmetry originates from the lack of helical edge states for $-k_{x2}\leq k_{x}\leq-k_{x1}$ and $k_{x1}\leq k_{x}\leq k_{x2}$, and it is caused by the specific spin configuration on the Fermi surface in PUMs. In $p_{x}$ and $p_{y}$-wave SC junctions, the tunneling conductance is symmetric for $eV$ [Fig. 6 (e) and Fig. 6(f)]. These line shapes are the same as those in normal metal/SC junctions shown in Fig. 6 (b) and Fig. 6(c): V-shape for the $p_{x}$-wave [Fig. 6 (e)] and ZBCP for the $p_{y}$-wave SC junctions [Fig. 6 (f)] [10, 11, 4]. Thus, the line shape of the tunneling conductance in the helical $p$-wave SC junction can be asymmetric for $eV$ in the presence of the PUM order, and in the spin-triplet $p_{x}$ and $p_{y}$-wave SC junctions, it is almost independent of the PUM order.
@@ III.3 Chiral $d$ and chiral $p$-wave superconductor junctions
In this subsection, we investigate the case of the chiral $d$ and chiral $p$-wave SC junctions without time-reversal symmetry. As well as the previous subsections, we first show the momentum-resolved tunneling conductance shown in Fig. 7. $\sigma(eV,k_{x})$ is enhanced at the corresponding chiral edge states inside the energy gap $-\Delta\leq eV\leq\Delta$ in the chiral $d$ and chiral $p$-wave superconductors [Figs. 7 (a)(b)(c)(d)]. Then $\sigma(eV,k_{x})$ in the chiral $d$-wave SC junctions at $t_{J}=0.25t$ and that of the chiral $p$-wave ones at $t_{J}=0,0.25t$ cross the zero bias voltage [Figs. 7 (b)(c)(d)(f)(g)(h)(j)(k)(l)], while $\sigma(eV,k_{x})$ in the chiral $d$-wave ones at $t_{J}=0$ does not cross $eV=0$ [Figs. 7 (a)(e)(i)]. It means that chiral edge states, which can contribute to the conductance at $eV=0$ appears in the chiral $d$-wave SC junctions in the presence of the PUM order. When we see $\sigma_{\uparrow}(eV,k_{x})/\sigma_{\mathrm{N}}$ ($\sigma_{\downarrow}(eV,k_{x})/\sigma_{\mathrm{N}}$), the range of $k_{x}$ contributing to the conductance is also restricted for $-k_{x2}\leq k_{x}\leq k_{x1}$ [Fig. 7 (f) and Fig. 7 (h)] ($-k_{x1}\leq k_{x}\leq k_{x2}$) [Fig. 7 (j) and Fig. 7 (l)]. We also find $\sigma_{\uparrow}(eV,k_{x})=\sigma_{\downarrow}(-eV,-k_{x})$.
@@
The tunneling conductance $\bar{\sigma}(eV)$ shown in Fig. 8 (c) and Fig. 8 (d) at $t_{J}=0.25t$ is symmetric for $eV$ in the presence of the PUM order. As compared with Fig. 8 (a) and Fig. 8 (b) at $t_{J}=0$, the tunneling conductance with up-spin $\bar{\sigma}_{\uparrow}(eV)$ and that with down-spin $\bar{\sigma}_{\downarrow}(eV)$ is not symmetric for $eV$ inside the energy gap $-\Delta\leq eV\leq\Delta$. Owing to the lack of $\sigma_{\uparrow}(eV,k_{x})$ for $k_{x1}\leq k_{x}\leq k_{x2}$ [Figs. 7 (f)(h)] and that of $\sigma_{\downarrow}(eV,k_{x})$ for $-k_{x2}\leq k_{x}\leq-k_{x1}$ [Figs. 7 (j)(l)], we find that the spin-resolved tunneling conductance has different values for the spin sectors inside the energy gap $-\Delta\leq eV\leq\Delta$. It can be applied to spin current in superconducting spintronics [44, 46, 45, 125, 126, 127]. In Fig. 8 (e) and Fig. 8 (f), we plot $\bar{\sigma}(eV)$ at $eV=0$ as a function of $t_{J}$. In general, the magnetic order suppresses the tunneling conductance at $eV=0$ as shown in chiral $p$-wave SC junctions [Fig. 8 (f)], owing to the suppression of the Andreev reflection. In chiral $d$-wave SC junctions, $\sigma(eV,k_{x})$ cross the zero bias voltage at a finite momentum for $t_{J}=0.25t$ [Fig. 7 (b)], while it does not for $t_{J}=0$ [Fig. 7 (a)]. Whether the chiral edge states cross at zero energy or not is determined by the size of the Fermi surface in the normal metal or the magnet. In the present study, the size of the Fermi surface in PUMs, that is, $k_{x2}$, becomes larger with the increase of $t_{J}$, as shown in Fig. 2. As a result, the tunneling conductance at $eV=0$ is enhanced with the increase of $t_{J}$ in chiral $d$-wave SC junctions [Fig. 8 (e)]. In the Appendix, we demonstrate the tunneling conductance for the chiral $p$-wave SC junctions with the in-plane d-vector. The direction of the d-vector does not affect the tunneling conductance in the chiral $p$-wave SC junctions.
@@ IV Conclusion
In this paper, we studied the tunneling conductance in superconducting junctions with PUMs using the effective PUM model in Ref. [74]. In spin-singlet and spin-triplet $p_{x}$ and $p_{y}$-wave SC junctions, the line shape of the tunneling conductance is not sensitive to the PUM order, even though the range of $k_{x}$ contributing to the spin-resolved conductance is restricted. The tunneling conductance in the helical $p$-wave SC junction becomes asymmetric with respect to $eV$ due to the lack of helical edge states contributing to the conductance. It is caused by the spin-split feature of the Fermi surface in PUMs. In chiral $d$ and chiral $p$-wave SC junctions, we found that the tunneling conductance is sensitive to the direction of the injected spin because of the time reversal symmetry breaking in SCs.
We have clarified that our obtained results based on the effective model Hamiltonian of PUM with the time-reversal symmetry breaking in PUM/SC junctions are qualitatively similar to those in Ref. [108], where the exchange coupling appearing in the Hamiltonian for PUMs is described by the odd function of Fermi momentum $\hat{H}_{\mathrm{PUM}}\propto k$ preserving the time-reversal symmetry. Likewise, we expect a similar behavior of the tunneling conductance in PUM/SC junctions along the $x$-direction to Ref. [108]. In the present calculation, we chose the chemical potential as $\mu_{p}=-3.85t$ and the Fermi surface shown in Fig. 1 (b) for PUMs. We confirmed that our results are generic by changing the chemical potential of PUMs and SCs. Hence, it indicates that the configuration of the electronic structure described by the simplified Hamiltonian $\hat{H}_{\mathrm{PUM}}\propto k_{x},k_{y}$ [108, 73] does not qualitatively deviate from that of the effective model with time-reversal symmetry breaking proposed in Ref. [74], within the calculations of the tunneling conductance.
For future perspective, the Josephson junctions with PUM without the time-reversal symmetry need to be discussed. In Ref. [82], the Josephson current is calculated by using the simplified PUM Hamiltonian with time reversal symmetry [108]. Indeed, the role of the time-reversal symmetry breaking of PUM should also be clarified in Josephson junctions with PUMs. Also, the analysis of the pair amplitude, including the odd-frequency pairing [17] is needed beyond previous studies based on the simplified Hamiltonian [82, 102].
@@ Acknowledgements.
Y. F. and K. Y. acknowledge financial support from the Sumitomo Foundation. Y. F. also acknowledges the numerical calculation support from Okayama University. Y. T. acknowledges financial support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grants Nos. 23K17668, 24K00583, 24K00556, and 24K00578). We thank S. Ikegaya, J. Cayao, and B. Lu for the valuable discussions.
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible, and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.

### v2

@@ III.2 Spin-triplet superconductor junctions with the time-reversal symmetry
Next, we demonstrate the tunneling conductance of the spin-triplet SC junctions with time-reversal symmetry. We plot $\sigma(eV,k_{x})$, $\sigma_{\uparrow}(eV,k_{x})$, and $\sigma_{\downarrow}(eV,k_{x})$ for helical $p$-wave, $p_{x}$-wave, and $p_{y}$-wave SC junctions in Figs. 5 (a)(d)(g), Figs. 5 (b)(e)(h), and Figs. 5 (c)(f)(i), respectively. In the case of the helical $p$-wave junction, we obtain the missing of the helical edge states from $k_{x2}\leq k_{x}\leq-k_{x1}$ and $k_{x1}\leq k_{x}\leq k_{x2}$ at $eV\sim 0.8\Delta$ as seen from $\sigma(eV,k_{x})$ plotted in Fig. 5 (a). For the $p_{x}$-wave SC junctions, because we consider the junction along the $y$-direction, we obtain the nodal structure in the conductance [Fig. 5 (b)] and ZBCP is absent. On the other hand, in the $p_{y}$-wave SC junctions, because we consider the junction along the $y$-direction, we obtain ZBCP caused by the zero-energy ABSs of the $p_{y}$-wave pairing [Fig. 5 (c)]. When we see the tunneling conductance with up-spin $\sigma_{\uparrow}(eV,k_{x})/\sigma_{\mathrm{N}}$, the range of $k_{x}$ contributing to the conductance is restricted for $k_{x2}\leq k_{x}\leq k_{x1}$ [Figs. 5 (d)(e)(f)]. On the other hand, for tunneling conductance with down spin, $\sigma_{\downarrow}(eV,k_{x})/\sigma_{\mathrm{N}}$ becomes nonzero for $k_{x1}\leq k_{x}\leq k_{x2}$ [Figs. 5 (g)(h)(i)]. We also find the relation: $\sigma_{\uparrow}(eV,k_{x})=\sigma_{\downarrow}(eV,-k_{x})$.
The resulting tunneling conductance $\bar{\sigma}(eV)$ is plotted in Fig. 6 with $t_{J}=0$ and $t_{J}=0.25t$ shown in Figs. 6 (a)(b)(c) and Figs. 6 (d)(e)(f). For the helical $p$-wave SC junctions, in the absence of the PUM order, the tunneling conductance is symmetric for $eV$ [Fig. 6 (a)]. However, because the momentum resolved tunneling conductance $\sigma(eV,k_{x})$ is asymmetric for $eV$ [Fig. 5 (a)], $\bar{\sigma}(eV)$ shown in green line of Fig. 6 (d) is not symmetric for $eV$. Then the value of $\bar{\sigma}(eV)-\bar{\sigma}(-eV)]/\sigma_{\mathrm{N}}$ becomes finite, caused by this asymmetry [brown dotted line in Fig. 6 (d)]. This asymmetry originates from the missing of helical edge states for $k_{x2}\leq k_{x}\leq-k_{x1}$ and $k_{x1}\leq k_{x}\leq k_{x2}$ at $eV\sim 0.8\Delta$ [Figs. 5 (d)(g)], and it is caused by the specific spin configuration on the Fermi surface in PUMs. In $p_{x}$ and $p_{y}$-wave SC junctions, the tunneling conductance is symmetric for $eV$ [Fig. 6 (e) and Fig. 6(f)]. These line shapes are the same as those in normal metal/SC junctions shown in Fig. 6 (b) and Fig. 6(c): V-shape for the $p_{x}$-wave [Fig. 6 (e)] and ZBCP for the $p_{y}$-wave SC junctions [Fig. 6 (f)] [10, 11, 4]. Thus, the line shape of the tunneling conductance in the helical $p$-wave SC junction can be asymmetric for $eV$ in the presence of the PUM order, and in the spin-triplet $p_{x}$ and $p_{y}$-wave SC junctions, it is almost independent of the PUM order.
@@ III.3 Chiral $d$ and chiral $p$-wave superconductor junctions
In this subsection, we investigate the case of the chiral $d$ and chiral $p$-wave SC junctions without time-reversal symmetry. As well as the previous subsections, we first show the momentum-resolved tunneling conductance shown in Fig. 7. $\sigma(eV,k_{x})$ is enhanced at the corresponding chiral edge states inside the energy gap $\Delta\leq eV\leq\Delta$ in the chiral $d$ and chiral $p$-wave superconductors [Figs. 7 (a)(b)(c)(d)]. Then $\sigma(eV,k_{x})$ in the chiral $d$-wave SC junctions at $t_{J}=0.25t$ and that of the chiral $p$-wave ones at $t_{J}=0,0.25t$ cross the zero bias voltage [Figs. 7 (b)(c)(d)(f)(g)(h)(j)(k)(l)], while $\sigma(eV,k_{x})$ in the chiral $d$-wave ones at $t_{J}=0$ does not cross $eV=0$ [Figs. 7 (a)(e)(i)]. It means that chiral edge states, which can contribute to the conductance at $eV=0$ appears in the chiral $d$-wave SC junctions in the presence of the PUM order. When we see $\sigma_{\uparrow}(eV,k_{x})/\sigma_{\mathrm{N}}$ ($\sigma_{\downarrow}(eV,k_{x})/\sigma_{\mathrm{N}}$), the range of $k_{x}$ contributing to the conductance is also restricted for $k_{x2}\leq k_{x}\leq k_{x1}$ [Fig. 7 (f) and Fig. 7 (h)] ($k_{x1}\leq k_{x}\leq k_{x2}$) [Fig. 7 (j) and Fig. 7 (l)]. We also find $\sigma_{\uparrow}(eV,k_{x})=\sigma_{\downarrow}(-eV,-k_{x})$.
@@
The tunneling conductance $\bar{\sigma}(eV)$ shown in Fig. 8 (c) and Fig. 8 (d) at $t_{J}=0.25t$ is symmetric for $eV$ in the presence of the PUM order. As compared with Fig. 8 (a) and Fig. 8 (b) at $t_{J}=0$, the tunneling conductance with up-spin $\bar{\sigma}_{\uparrow}(eV)$ and that with down-spin $\bar{\sigma}_{\downarrow}(eV)$ is not symmetric for $eV$ inside the energy gap $\Delta\leq eV\leq\Delta$. Owing to the missing of $\sigma_{\uparrow}(eV,k_{x})$ for $k_{x1}\leq k_{x}\leq k_{x2}$ [Figs. 7 (f)(h)] and that of $\sigma_{\downarrow}(eV,k_{x})$ for $k_{x2}\leq k_{x}\leq-k_{x1}$ [Figs. 7 (j)(l)], we find that the spin-resolved tunneling conductance has different values for the spin sectors inside the energy gap $\Delta\leq eV\leq\Delta$. It can be applied to spin current in superconducting spintronics [44, 46, 45, 127, 128, 129]. In Fig. 8 (e) and Fig. 8 (f), we plot $\bar{\sigma}(eV)$ at $eV=0$ as a function of $t_{J}$. In general, the magnetic order suppresses the tunneling conductance at $eV=0$ as shown in chiral $p$-wave SC junctions [Fig. 8 (f)], owing to the suppression of the Andreev reflection. In chiral $d$-wave SC junctions, $\sigma(eV,k_{x})$ cross the zero bias voltage at a finite momentum for $t_{J}=0.25t$ [Fig. 7 (b)], while it does not for $t_{J}=0$ [Fig. 7 (a)]. Whether the chiral edge states cross at zero energy or not is determined by the size of the Fermi surface in the normal metal or the magnet. In the present study, the size of the Fermi surface in PUMs, that is, $k_{x2}$, becomes larger with the increase of $t_{J}$, as shown in Fig. 2. As a result, the tunneling conductance at $eV=0$ is enhanced with the increase of $t_{J}$ in chiral $d$-wave SC junctions [Fig. 8 (e)]. In the Appendix, we demonstrate the tunneling conductance for the chiral $p$-wave SC junctions with the in-plane d-vector. The direction of the d-vector does not affect the tunneling conductance in the chiral $p$-wave SC junctions.
@@ IV Conclusion
In this paper, we studied the tunneling conductance in superconducting junctions with PUMs using the effective PUM model in Ref. [75]. In spin-singlet and spin-triplet $p_{x}$ and $p_{y}$-wave SC junctions, the line shape of the tunneling conductance is not sensitive to the PUM order, even though the range of $k_{x}$ contributing to the spin-resolved conductance is restricted. The tunneling conductance in the helical $p$-wave SC junction becomes asymmetric with respect to $eV$ due to the missing of helical edge states contributing to the conductance. It is caused by the spin-split feature of the Fermi surface in PUMs. In chiral $d$ and chiral $p$-wave SC junctions, we found that the tunneling conductance is sensitive to the direction of the injected spin because of the time reversal symmetry breaking in SCs.
We have clarified that our obtained results based on the effective model Hamiltonian of PUM with the time-reversal symmetry breaking in PUM/SC junctions are qualitatively similar to those in Ref. [110], where the exchange coupling appearing in the Hamiltonian for PUMs is described by the odd function of Fermi momentum $\hat{H}_{\mathrm{PUM}}\propto k$ preserving the time-reversal symmetry. In our results, we chose the junction along the $y$-direction, where the lattice translation symmetry is preserved. As a result, our calculations do not change as compared to the simplified continuum model [110]. In the present calculation, we chose the chemical potential as $\mu_{p}=-3.85t$ and the Fermi surface shown in Fig. 1 (b) for PUMs. We confirmed that our results are generic by changing the chemical potential of PUMs and SCs. Indeed, the configuration of the electronic structure described by the simplified Hamiltonian $\hat{H}_{\mathrm{PUM}}\propto k_{y}$ [110, 74] does not qualitatively deviate from that of the effective model with time-reversal symmetry breaking proposed in Ref. [75], within the calculations of the tunneling conductance. Therefore, our results are important for the justification of the phenomenological analysis in the simplified continuum model in Ref. [110]. While, along the $x$-direction, because the system does not have the translational symmetry along the direction which is normal to the interface, $C_{2}||t]-$symmetry [75] is broken. Thus, the effect of the time-reversal symmetry breaking can be expected in junctions along the $x$-direction.
@@
In our work, we set the isotropic tunneling Hamiltonian at the interface. For future perspective, to enhance the effect of time-reversal symmetry breaking, we can focus on the anisotropic tunneling hopping at the interface and another direction of junctions ($x$-direction). In addition, the Josephson junctions with PUM without time-reversal symmetry need to be discussed. In Ref. [84], the Josephson current is calculated by using the simplified PUM Hamiltonian with time reversal symmetry [110]. Indeed, the role of time-reversal symmetry breaking in PUM should also be clarified in Josephson junctions with PUMs. Also, the analysis of the pair amplitude, including the odd-frequency pairing [17], is needed beyond previous studies based on the simplified Hamiltonian [84, 104].
@@ Acknowledgements.
Y. F. and K. Y. acknowledge financial support from the Sumitomo Foundation. Y. F. also acknowledges the numerical calculation support from Okayama University. Y. T. acknowledges financial support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grants Nos. 23K17668, 24K00583, 24K00556, and 24K00578). We thank S. Ikegaya, J. Cayao, B. Lu, and P. Sukhachov for the valuable discussions.